OpenStudy (anonymous):

Two cars raced at a race track. The faster car traveled 20 mph faster than the slower car. In the time that the slower car traveled 165 miles, the faster car traveled 225 miles. If the speeds of the cars remained constant, how fast did the slower car travel during the race?

OpenStudy (anonymous):

options are 1. 15 mph 2. 55 mph 3. 75 mph 4. 130 mph

OpenStudy (***[isuru]***):

hi, if the speed of slower car is "x" then the speed of faster car is (20 + x) so both cars traveled 165 miles and 225 miles in the same time. Let's assume that time period as "t".Remember that "t" is common for the motion of both cars... for the slower car... distance traveled = speed * duration 165 = xt ------(1) for the faster car ... distance traveled = speed * duration 225 = t(x+20) ----(2) divide (1) by (2) 165/225 = x/(x+20) 11/15 = x/(x+20) 11x + 210 = 15x 4x = 220 x = 55 mph hope this will help ya!

OpenStudy (anonymous):

55 mph

OpenStudy (anonymous):

thanks jake and isuru

OpenStudy (***[isuru]***):

u r welcome!!